# Segment Any 3D Gaussians

### v2

@@ A.4 SAGA with Hybrid / Implicit Radiance Fields
Although SAGA is designed for segmentation in 3D-GS, its scale gate mechanism is not confined to a particular type of radiance field. We demonstrate this versatility by adapting SAGA to GARField, replacing the scale-conditioned affinity field with a scale-gated affinity field. As illustrated in Figure 5, the performance remains competitive.
Figure 5: Adapting the scale gate mechanism with GARField achieves competitive results, demonstrating the potential of SAGA across different radiance fields.
A.5 More Qualitative Results
We present additional qualitative results in Figure 6 and Figure 7.
Figure 6: More qualitative results of SAGA.
Figure 7: More qualitative results of SAGA.

### v3

@@ A.5 SAGA with Hybrid Radiance Fields
Although SAGA is designed for segmentation in 3D-GS, its scale gate mechanism is not confined to a particular type of radiance field. We demonstrate this versatility by adapting SAGA to GARField, replacing the scale-conditioned affinity field with a scale-gated affinity field. As illustrated in Figure A1, the performance remains competitive.
Furthermore, we apply SAGA to InstantNGP (Müller et al. 2022), which can be viewed as an explicit version of GARField without the MLP. As shown in Figure A2, SAGA continues to produce competitive results. It is worth noting that direct clustering on the hash grid of InstantNGP is infeasible. To address this, we follow GARField’s approach, using 3D-GS to query the hash grids and extract 3D features for clustering.
Figure A2: Applying SAGA to Instant-NGP achieves competitive segmentation performance, further demonstrating the generalizability and robustness of SAGA across different radiance field representations.
A.6 More Qualitative Results
We present additional qualitative results in Figures A5 and A4. In Figure A3, it is evident that the affinity features in SAGA exhibit better stability compared to GARField. We perform clustering on the affinity features across the entire scene. At a smaller scale (0.01), both SAGA and GARField achieve fine-grained segmentation. However, when conducting coarse segmentation at a larger scale, GARField tends to merge smaller objects into larger ones (e.g., the table). This behavior can be attributed to two factors: first, GARField assumes that the entire scene belongs to the same “object” when the scale exceeds that of any existing objects within the scene. Second, GARField employs an implicit feature field to fit the scale-conditioned affinity features. As noted by Mildenhall et al. (2020), radiance fields often produce over-smoothed predictions, which can result in the loss of small objects at larger scales. In contrast, SAGA assigns an explicit affinity feature to each 3D Gaussian in the 3D-GS model and uses a simple one-layer linear layer (i.e., the scale gate) to model the scale-conditioned effect. This approach allows SAGA to better preserve smaller objects. This is also evidenced by Figure A1: by replacing the multi-layer perceptron of GARField with our proposed scale-gate mechanism, GARField can preserve the small objects on the table.
Figure A3: Qualitative comparison with GARField. We conduct feature clustering across the whole scene (LERF-figurines). Compared to GARField, which employs an additional feature field to model affinity features, SAGA demonstrates greater stability by utilizing explicit affinity features. At larger scales, SAGA effectively preserves the perception of small objects without merging them with other targets.
Figure A4: More qualitative results of SAGA.
Figure A5: More qualitative results of SAGA.